OpenStudy (anonymous):

how do you find the polar equation of the parabola with the given vertex (2, pi/2) and focus at the pole? I AM SO LOST HELP!!!

OpenStudy (anonymous):

but i want to find a polar equaiton

OpenStudy (anonymous):

someone help me.....

OpenStudy (anonymous):

focus at the pole?

OpenStudy (unklerhaukus):

oh, 

OpenStudy (unklerhaukus):

circle \[x^2+y^2=r^2\]\[y^2=r^2-x^2\] parabola \[y=-x^2+a\]\[y+x^2=a\]

OpenStudy (anonymous):

yes focus at the pole and they said that the directrix is 4 which doesnt make any sense...how did they get the directrix 4? and the ecctricity 1..?

OpenStudy (unklerhaukus):